# RGFN: Synthesizable Molecular Generation Using GFlowNets

### v1

@@ 1 Introduction
We experimentally evaluate RGFN on a set of diverse screening tasks, including docking score approximation with a trained proxy model for soluble epoxide hydrolase (sEH), GPU-accelerated direct docking score calculations for multiple protein targets (Mpro, ClpP, TBLR1 and sEH), and biological activity estimation with a trained proxy model for senolytic classification [64]. We demonstrate that RGFN produces similar optimization quality and diversity to existing fragment-based approaches while ensuring straightforward synthetic routes for predicted hit compounds.
@@ 2 Related work
Generative models for molecular discovery. A plethora of methods have been developed for molecular generation [42, 7] using machine learning. These methods can be categorized depending on the molecular representation used, including textual representations such as SMILES [30, 3, 34], molecular graphs [29, 40, 46] or 3D atom coordinate representations [45], as well as the underlying methodology, for example, variational autoencoders [29, 40], reinforcement learning [46, 33] or diffusion models [51]. Recently, Generative Flow Networks (GFlowNets) [4, 43, 50, 54, 62] have emerged as a promising paradigm for molecular generation due to their ability to sample large and diverse candidate small molecule space, which is crucial in the drug discovery process. Traditionally, GFlowNets operated on the graph representation level, and candidate molecules were generated as a sequence of actions in which either individual atoms or small molecular fragments were combined to form a final molecule. While using graph representations, as opposed to textual or 3D representations, allows the enforcement of the validity of the generated molecules, it does not guarantee a valid route for chemical synthesis. Here, we expand on the GFlowNet framework by modifying the space of actions to consist of choosing molecular fragments and executing compatible chemical reactions/transformations, in turn guaranteeing both physical-chemical validity and synthesizability.
Synthesizability in generative models. One approach to ensuring the synthesizability of generated molecules is by using a scoring function, either utilizing it as one of the optimization criteria [33], or as a post-processing step for filtering generated molecules. Multiple scoring approaches, both heuristic [16, 20] and ML-based [37], exist in the literature. However, synthesizability estimation is difficult in practice, and can fail to generalize out-of-distribution in the case of ML models. Furthermore, theoretical synthesizability does not necessarily account for the cost of synthesis. Because of this, a preferable approach might be to constrain the space of possible molecules to those easily synthesized by operating in a predefined space of chemical reactions and fragments. Several recent strategies employ this approach [8, 19, 57], including reinforcement learning-based methods [21, 24]. We extend this line of investigation not only by translating the concept to the GFlowNet framework but also by proposing a curated set of robust chemical reactions and fragments that ensure efficient synthesis at low total costs.
@@ 3.2 Reaction-GFlowNet
$p(m^{\prime}_{i}|r)=\sigma^{|M^{\prime}|}(\mathbf{s})_{i},\;s_{i}=\mathrm{MLP}_{M^{\prime}}(f(m^{\prime}_{i},r)),$ (5)
@@ 3.3 Chemical language
We select seventeen reactions and 350 building blocks for our model. These include amide bond formation, nucleophilic aromatic substitution, Michael addition, isocyanate-based urea synthesis, sulfur fluoride exchange (SuFEx), sulfonyl chloride substitution, alkyne-azide and nitrile-azide cycloadditions, esterification reactions, urea synthesis using carbonyl surrogates, Suzuki-Miyaura, Buchwald-Hartwig, and Sonogashira cross-couplings, amide reduction, and peptide terminal thiourea cyclization reactions to produce iminohydantoins and tetrazoles. The chosen reactions are known to be typically quite robust [9] and are often high-yielding (75-100%), thus enforcing reliable synthesis pathways when sampling molecules from our model. To simulate couplings in Python, reactions are encoded as SMARTS templates. To ensure compatible building blocks yielding specific, chemically valid products and to enable parent state computation, we introduce multiple variants corresponding to differing reagent types for most of the proposed reactions. In some cases SMARTS templates encode reactions where one of the reagents is not specified. We describe these transformations as implicit reactions (Appendix D). Additionally, we introduce surrogate reactions where the SMARTS templates and SMILES strings encode for alternate building blocks (see Section C.2 for more details). Finally, once again for the sake of specificity, reactions are duplicated by swapping the order of the building block reactants. In total, 132 different SMARTS templates are used.
@@
During the construction of the curated building block database, only affordable reagents (i.e., building blocks) are considered. For the purposes of this study we define affordable reagents to be those priced at less than or equal to \$200 per gram. The mean cost per gram of reagents selected for this study is \$22.52, the lowest cost \$0.023 per gram, and the highest cost \$190 per gram (see Appendix M for more details on cost estimation).
@@
A crucial consideration when choosing the set of reactions and fragments used is the state space size. This is difficult to compute precisely since a different set of reactions or building blocks is valid for every state in a given trajectory. We estimate this based on 1,000 random trajectories instead (details can be found in Appendix A). In addition to our 350 low-cost fragments, we also perform this analysis with 8,000 additional random Enamine building blocks. Comparison for different numbers of maximum reactions is presented in Figure 2. We demonstrate that even with curated low-cost reactants and limiting the number to a maximum of four reactions, state space size is an order of magnitude greater than the number of molecules contained in Enamine REAL [15]. This size can increase significantly with the addition of more fragments and/or an increase in the maximum number of reactions. Additional discussions regarding scaling can be found in Section 4.3.
@@ 4.1 Set-up
Throughout the course of the conducted computational experimental study, we aim to evaluate the performance of the proposed approach across several diverse biological oracles of interest. This includes proxy models (machine learning oracles, pretrained on the existing data and used for higher computational efficiency): first, the commonly used sEH proxy as described in [4]. Second, a graph neural network trained on the biological activity classification task of senolytic [64] recognition. Proxy model details are provided in Section B.1.
Per the GFlowNet training algorithm, the reward is calculated for a batch of dozens to hundreds of molecules at each training step, rendering traditional computational docking score algorithms like AutoDock Vina [60] infeasible for very large training runs. As a result, previous applications of GFlowNets to biological design [4, 54] employed a fast pre-trained proxy model trained on docking scores instead. These proxies, while lightweight, present potential issues should the GFlowNet generate molecules outside their training data distributions and require receptor-specific datasets. To circumvent this, we use the GPU-accelerated Vina-GPU 2.1 [58] implementation of the QuickVina 2 [2] docking algorithm to calculate docking scores directly in the training loop of RGFN. This approach allows for drastically increased flexibility in protein target selection while eliminating proxy generalization failure. We selected X-ray crystal structures of human soluble epoxy hydrolase (sEH), ATP-dependent Clp protease proteolytic subunit (ClpP), SARS-CoV-2 main protease (Mpro), and transducin $\beta$-like-related protein 1 as targets for evaluating RGFN using a docking reward (detailed motivation for specific target selection is provided in Appendix F).
@@ 4.2 Comparison with existing methods
We begin experimental evaluation with a comparison to several state-of-the-art methods for molecular discovery. Specifically, we consider a genetic algorithm operating on molecular graphs (GraphGA) [28] as implemented in [10], which has been demonstrated to be a very strong baseline for molecular discovery [18], Monte Carlo tree search-based SyntheMol [57], and a fragment-based GFlowNet (FGFN) [4] as implemented in [48]. Training details can be found in Section B.2. It is worth noting that besides SyntheMol, which also operates in the space of chemical reactions and building blocks derived from the Enamine database, our remaining benchmarks do not explicitly enforce synthesizability when generating molecules. Because of this, in this section, we will examine not only the quality of generated molecules in terms of optimized properties but also their synthesizability. We consider only a single reaction-based approach, as other existing methods employing this paradigm [24, 21] do not share code or curated reactions and building blocks, making reproduction difficult.
We first examine the distributions of rewards found by each method across three different oracles used for training: sEH proxy, senolytic proxy, and GPU-accelerated docking for ClpP. The results are presented in Figure 3. As can be seen, while RGFN underperforms in terms of average reward when compared to the method not enforcing synthesizability (GraphGA), it outperforms SyntheMol’s reaction-based sampling. Interestingly, when compared to standard FGFN, RGFN either performs similarly (ClpP docking) or achieves higher average rewards. This is most striking in the case of the challenging senolytic discovery task, in which a proxy is trained on a severely imbalanced dataset with less than 100 actives, resulting in a sparse reward function. We suspect that this, possibly combined with a lack of compatibility between the FGFN fragments and known senolytics, led to the failure to discover any high-reward molecules. However, RGFN succeeds in the task and finds a wide range of senolytic candidates.
@@
Secondly, we examine the number of discovered modes for each method, with a mode defined as a molecule with computed reward above a threshold (sEH: 7, senolytics: 50, ClpP docking: 10), and Tanimoto similarity to every other mode $<0.5$. The number of discovered modes across tasks as a function of normalized iterations is presented in Figure 4. Note that in the case of GraphGA, FGFN, and RGFN this simply translates to the number of oracle calls, but for SyntheMol, due to large computational overhead, we impose a maximum number of rollouts such that training time was comparable to RGFN (see Section B.2 for details). As can be seen, despite slightly worse average rewards, FGFN still outperforms other methods in terms of the number of discovered modes (with the exception of senolytic discovery task, where it fails to discover any high-reward molecules). This suggests that RGFN samples are less diverse, possibly due to the relatively small number of fragments and reactions used. However, RGFN still outperforms remaining methods across all tasks, suggesting that it preserves some of the benefits of the diversity-focused GFlowNet framework.
@@
Finally, we evaluate the synthesizability of the generated compounds as a key output. We present average values of several synthesizability-related metrics, computed over top-k modes generated for each method, in Table 1. We include measures indicating average molecular weight and drug-likeness (QED) to gauge the size of generated compounds. Furthermore, for completeness, we also include SAScores [16], but note that they are only a rough approximation of ease of synthesis. For a better estimate of synthesizability we perform retrosynthesis using AiZynthFinder [20] and count the average number of molecules for which a valid retrosynthesis pathway was found. Note that to reduce variance, we compute molecular weight, QED, and SAScores over the top-500 modes, but due to high computational cost, AiZynthFinder scores are computed only over top-100 modes. As can be seen, while there is some variance across tasks, RGFN performs similarly to SyntheMol in terms of both synthesizability scores, and significantly outperforms GraphGA and FGFN. All RGFN modes were additionally inspected manually by an expert chemist and confirmed as synthesizable, which indicates that AiZynth scores are likely underestimated.
@@ 4.3 Scaling to larger sets of fragments
Next we investigate the influence of a fragment embedding scheme proposed in Section 3.2. In the standard implementation of the GFlowNet policy, actions are represented as independent embeddings in the MLP. These encode actions as indices, effectively disregarding their respective structures and all information contained therein. The model must thus select from a library of reagents without any knowledge as to their chemical makeup or properties. While finding similarities between actions may be a relatively easy task for small action spaces, it becomes more difficult when the size of the action space increases. To scale RGFN to a larger size of the building block library, we proposed to encode building block selection actions using molecular fingerprints, allowing the model to leverage their internal structures without any additional computational overhead during inference. In Figure 5, we observe that our fingerprint embedding scheme allows for drastically faster convergence compared to the standard independent action embedding, especially for large library sizes. The details on how the larger fragment libraries were created can be found in Appendix E.

### v2

@@ Abstract
00footnotetext: Source code available at https://github.com/koziarskilab/RGFN.
@@ 1 Introduction
We experimentally evaluate RGFN on a set of diverse screening tasks, including docking score approximation with a trained proxy model for soluble epoxide hydrolase (sEH), GPU-accelerated direct docking score calculations for multiple protein targets (Mpro, ClpP, TBLR1 and sEH), and biological activity estimation with a trained proxy model for dopamine receptor type 2 (DRD2) receptor activity and senolytic activity [74]. We demonstrate that RGFN produces similar optimization quality and diversity to existing fragment-based approaches while ensuring straightforward synthetic routes for predicted hit compounds.
@@ 2 Related work
Generative models for molecular discovery. A plethora of methods have been developed for molecular generation [48, 7] using machine learning. These methods can be categorized depending on the molecular representation used, including textual representations such as SMILES [34, 3, 39], molecular graphs [33, 46, 54] or 3D atom coordinate representations [52], as well as the underlying methodology, for example, variational autoencoders [33, 46], reinforcement learning [54, 37] or diffusion models [60]. Recently, Generative Flow Networks (GFlowNets) [4, 50, 59, 64, 72, 19] have emerged as a promising paradigm for molecular generation due to their ability to sample large and diverse candidate small molecule space, which is crucial in the drug discovery process. Traditionally, GFlowNets for molecular generation operated on the graph representation level, and candidate molecules were generated as a sequence of actions in which either individual atoms or small molecular fragments were combined to form a final molecule. While using graph representations, as opposed to textual or 3D representations, allows the enforcement of the validity of the generated molecules, it does not guarantee a straightforward route for chemical synthesis. Here, we expand on the GFlowNet framework by modifying the space of actions to consist of choosing molecular fragments and executing compatible chemical reactions/transformations, in turn guaranteeing both physical-chemical validity and synthesizability.
Synthesizability in generative models. One approach to ensuring the synthesizability of generated molecules is by using a scoring function, either utilizing it as one of the optimization criteria [37], or as a post-processing step for filtering generated molecules. Multiple scoring approaches, both heuristic [18, 23] and ML-based [42], exist in the literature. Another branch of research focuses on using reaction models and traversing predicted synthesis graph [8, 12, 38, 55]. However, reaction and synthesizability estimation is difficult in practice, and can fail to generalize out-of-distribution in the case of ML models. Furthermore, theoretical synthesizability does not necessarily account for the cost of synthesis. Because of this, a preferable approach might be to constrain the space of possible molecules to those easily synthesized by operating in a predefined space of chemical reactions and fragments. Several recent strategies employ this approach [22, 49, 67], including reinforcement learning-based methods [24, 27] and a concurrent work utilizing GFlowNets [15]. We extend this line of investigation not only by translating the concept to the GFlowNet framework but also by proposing a curated set of robust chemical reactions and fragments that ensure efficient synthesis at low total costs.
@@ 3.1 Generative Flow Networks
Another way to rephrase the flow-matching constraints is to learn a forward policy $P_{F}(s_{i+1}|s_{i})$ such that trajectories starting at $s_{0}$ and taking actions sampled by $P_{F}$ terminate at $x\in\mathcal{X}$ proportional to the reward.
Trajectory balance. Several training losses have been explored to train GFlowNets. Among these, trajectory balance [45] has been shown to improve credit assignment. In addition to learning a forward policy $P_{F}$, we also learn a backward policy $P_{B}$ and a scalar $Z_{\theta}$, such that, for every trajectory $\tau=(s_{0}\rightarrow s_{1}\rightarrow\dots\rightarrow s_{n}=x)$, they satisfy:
$Z_{\theta}\prod_{t=1}^{n}P_{F}(s_{t}|s_{t-1})=R(x)\prod_{t=1}^{n}P_{B}(s_{t-1}|s_{t})$ (2)
@@ 3.2 Reaction-GFlowNet
$p(m^{\prime}_{i})=\sigma^{|M^{\prime}|}(\mathbf{s})_{i},\;s_{i}=\mathrm{MLP}_{M^{\prime}}(f(m^{\prime}_{i})),$ (6)
@@ 3.3 Chemical language
We select seventeen reactions and 350 building blocks for our model. These include amide bond formation, nucleophilic aromatic substitution, Michael addition, isocyanate-based urea synthesis, sulfur fluoride exchange (SuFEx), sulfonyl chloride substitution, alkyne-azide and nitrile-azide cycloadditions, esterification reactions, urea synthesis using carbonyl surrogates, Suzuki-Miyaura, Buchwald-Hartwig, and Sonogashira cross-couplings, amide reduction, and peptide terminal thiourea cyclization reactions to produce iminohydantoins and tetrazoles. The chosen reactions are known to be typically quite robust [9] and are often high-yielding (75-100%), thus enforcing reliable synthesis pathways when sampling molecules from our model. To simulate couplings in Python, reactions are encoded as SMARTS templates. To ensure compatible building blocks yielding specific, chemically valid products and to enable parent state computation, we introduce multiple variants corresponding to differing reagent types for most of the proposed reactions. In some cases SMARTS templates encode reactions where one of the reagents is not specified. We describe these transformations as implicit reactions (Appendix E). Additionally, we introduce surrogate reactions where the SMARTS templates and SMILES strings encode for alternate building blocks (see Section C.2 for more details). Finally, once again for the sake of specificity, reactions are duplicated by swapping the order of the building block reactants. In total, 132 different SMARTS templates are used.
@@
During the construction of the curated building block database, only affordable reagents (i.e., building blocks) are considered. For the purposes of this study we define affordable reagents to be those priced at less than or equal to \$200 per gram. The mean cost per gram of reagents selected for this study is \$22.52, the lowest cost \$0.023 per gram, and the highest cost \$190 per gram (see Appendix N for more details on cost estimation).
@@
A crucial consideration when choosing the set of reactions and fragments used is the state space size (the number of possible molecules that can be generated using our framework). This is difficult to compute precisely since a different set of reactions or building blocks is valid for every state in a given trajectory. We estimate this based on 1,000 random trajectories instead (details can be found in Appendix A). In addition to our 350 low-cost fragments, we also perform this analysis with 8,000 additional random Enamine building blocks. Comparison for different numbers of maximum reactions is presented in Figure 2. We demonstrate that even with curated low-cost reactants and limiting the number to a maximum of four reactions, state space size is an order of magnitude greater than the number of molecules contained in Enamine REAL [17]. This size can increase significantly with the addition of more fragments and/or an increase in the maximum number of reactions. Additional discussions regarding scaling can be found in Section 4.3.
@@ 4.1 Set-up
Throughout the course of the conducted computational experimental study, we aim to evaluate the performance of the proposed approach across several diverse biological oracles of interest. This includes proxy models (machine learning oracles, pretrained on the existing data and used for higher computational efficiency): first, the commonly used sEH proxy as described in [4]. Second, a graph neural network trained on the biological activity classification task of senolytic [74] recognition. Third, the Dopamine Receptor D2 (DRD2) oracle [53] from Therapeutics Data Commons [28]. Proxy model details are provided in Section B.1.
Per the GFlowNet training algorithm, the reward is calculated for a batch of dozens to hundreds of molecules at each training step, rendering traditional computational docking score algorithms like AutoDock Vina [70] infeasible for very large training runs. As a result, previous applications of GFlowNets to biological design [4, 64] employed a fast pre-trained proxy model trained on docking scores instead. These proxies, while lightweight, present potential issues should the GFlowNet generate molecules outside their training data distributions and require receptor-specific datasets. To circumvent this, we use the GPU-accelerated Vina-GPU 2.1 [68] implementation of the QuickVina 2 [2] docking algorithm to calculate docking scores directly in the training loop of RGFN. This approach allows for drastically increased flexibility in protein target selection while eliminating proxy generalization failure. We selected X-ray crystal structures of human soluble epoxy hydrolase (sEH), ATP-dependent Clp protease proteolytic subunit (ClpP), SARS-CoV-2 main protease (Mpro), and transducin $\beta$-like-related protein 1 as targets for evaluating RGFN using a docking reward (detailed motivation for specific target selection is provided in Appendix G).
@@ 4.2 Comparison with existing methods
We begin experimental evaluation with a comparison to several state-of-the-art methods for molecular discovery. Specifically, we consider a genetic algorithm operating on molecular graphs (GraphGA) [32] as implemented in [10], which has been demonstrated to be a very strong baseline for molecular discovery [21], Monte Carlo tree search-based SyntheMol [67], cascade variational autoencoder (casVAE) [49], and a fragment-based GFlowNet (FGFN) [4] as implemented in [57]. For FGFN, we additionally considered its variant that had a SAScore as one of the reward terms (FGFN+SA). Training details can be found in Section B.2. It is worth noting that besides SyntheMol, which also operates in the space of chemical reactions and building blocks derived from the Enamine database, and casVAE, which used the set of reaction trees obtained from the USPTO database [43], our remaining benchmarks do not explicitly enforce synthesizability when generating molecules. Because of this, in this section, we will examine not only the quality of generated molecules in terms of optimized properties but also their synthesizability. We consider only two reaction-based approaches, as other existing methods employing this paradigm [27, 24] do not share code or curated reactions and building blocks, making reproduction difficult.
We first examine the distributions of rewards found by each method across four different oracles used for training: sEH proxy, senolytic proxy, DRD2 proxy, and GPU-accelerated docking for ClpP. The results are presented in Figure 3. As can be seen, while RGFN underperforms in terms of average reward when compared to the method not enforcing synthesizability (GraphGA), it outperforms SyntheMol’s and casVAE’s reaction-based sampling. Interestingly, when compared to standard FGFN, RGFN either performs similarly (ClpP docking) or achieves higher average rewards. This is most striking in the case of the challenging senolytic discovery task, in which a proxy is trained on a severely imbalanced dataset with less than 100 actives, resulting in a sparse reward function. We suspect that this, possibly combined with a lack of compatibility between the FGFN fragments and known senolytics, led to the failure to discover any high-reward molecules. However, RGFN succeeds in the task and finds a wide range of senolytic candidates. Finally, the gap in performance is even larger between RGFN and FGFN+SA, indicating that introducing synthesizability constraints reduces the ability of FGFN to discover high-reward molecules.
@@
Secondly, we examine the number of discovered modes for each method, with a mode defined as a molecule with computed reward above a threshold (sEH: 7, senolytics: 50, DRD2: 0.95, ClpP docking: 10), and Tanimoto similarity to every other mode $<0.5$. We use Leader algorithm for mode computation. The number of discovered modes across tasks as a function of normalized iterations is presented in Figure 4. Note that in the case of GraphGA, FGFN, FGFN+SA, and RGFN this simply translates to the number of oracle calls, but for SyntheMol and casVAE, due to large computational overhead, we impose a maximum number of oracle calls such that training time was comparable to RGFN (see Section B.2 for details). Note that in the case of casVAE, this resulted in a very small number of molecules being visited in the allotted time. As can be seen, despite slightly worse average rewards, FGFN still outperforms other methods in terms of the number of discovered modes (with the exception of senolytic discovery task, where it fails to discover any high-reward molecules). This includes FGFN+SA, despite its generally worse performance than FGFN. This suggests that RGFN samples are less diverse, possibly due to the relatively small number of fragments and reactions used. However, RGFN still outperforms remaining methods across all tasks, suggesting that it preserves some of the benefits of the diversity-focused GFlowNet framework.
@@
Finally, we evaluate the synthesizability of the generated compounds as a key output. We present average values of several synthesizability-related metrics, computed over top-k modes generated for each method, in Table 1. We include measures indicating average molecular weight and drug-likeness (QED) to gauge the size of generated compounds. Furthermore, for completeness, we also include SAScores [18], but note that they are only a rough approximation of ease of synthesis. For a better estimate of synthesizability we perform retrosynthesis using AiZynthFinder [23] and count the average number of molecules for which a valid retrosynthesis pathway was found. However, it is important to note that both SAScores and AiZynthFinder scores are inherently noisy metrics. While we use them to evaluate all methods for the sake of rigorousness, ultimately molecules generated by RGFN (as well as SyntheMol and casVAE) are highly likely to be synthesizable. Note that to reduce variance, we compute molecular weight, QED, and SAScores over the top-500 modes, but due to high computational cost, AiZynthFinder scores are computed only over top-100 modes. As can be seen, while there is some variance across tasks, RGFN performs similarly to SyntheMol and casVAE in terms of both synthesizability scores, and significantly outperforms GraphGA and FGFN. Crucially, including SAScore as a reward does improve the performance of FGFN in terms of that metric, but does not drastically change the AiZynthFinder scores, demonstrating that it is insufficient to guarantee synthesizability. All RGFN modes were additionally inspected manually by an expert chemist and confirmed as synthesizable, which indicates that AiZynth scores are likely underestimated.
@@
DRD2 GraphGA 475.4 ± 53.2 0.42 ± 0.12 2.50 ± 0.23 0.41
SyntheMol 365.6 ± 54.3 0.72 ± 0.14 2.78 ± 0.43 0.66
casVAE 404.8 ± 83.5 0.59 ± 0.20 2.42 ± 0.38 0.87
FGFN 386.5 ± 45.0 0.63 ± 0.11 2.58 ± 0.54 0.76
FGFN+SA 381.1 ± 35.1 0.64 ± 0.10 2.37 ± 0.37 0.78
RGFN 447.1 ± 45.7 0.44 ± 0.10 2.79 ± 0.34 0.87
@@ 4.3 Scaling to larger sets of fragments
Next we investigate the influence of a fragment embedding scheme proposed in Section 3.2. In the standard implementation of the GFlowNet policy, actions are represented as independent embeddings in the MLP. These encode actions as indices, effectively disregarding their respective structures and all information contained therein. The model must thus select from a library of reagents without any knowledge as to their chemical makeup or properties. While finding similarities between actions may be a relatively easy task for small action spaces, it becomes more difficult when the size of the action space increases. To scale RGFN to a larger size of the building block library, we proposed to encode building block selection actions using molecular fingerprints, allowing the model to leverage their internal structures without any additional computational overhead during inference. In Figure 5, we observe that our fingerprint embedding scheme allows for drastically faster convergence compared to the standard independent action embedding, especially for large library sizes. The details on how the larger fragment libraries were created can be found in Appendix F.